# Multi-orbital two-particle self-consistent approach – strengths and limitations

### v1

@@ III.3 Numerical benchmarks
Keeping the above described shortcomings in mind, we are now exploring the range of applicability of multi-orbital TPSC by testing the reliability of TPSC3 and TPSC5 through comparison to more accurate numerical approaches. In a first step, we exclusively aim at validating the quality of the predicted TPEVs, as these form the basis for the TPSC approach. For this we compare to both ED and DMFT.
III.3.1 Comparison to ED
First, we compare the predicted TPEV between ED an TPSC, for a two-orbital dimer with Hubbard-Kanamori interactions
where we introduced a Hubbard-Kanamori type interaction given as
$H^{HK}$ $=U\hat{n}_{o_{1}^{\prime}}^{\uparrow}\hat{n}_{o_{1}^{\prime}}^{\downarrow}+\sum_{{o_{1}^{\prime}}\neq{o_{2}^{\prime}}}\left((U-2J)\hat{n}_{o_{1}^{\prime}}^{\uparrow}\hat{n}_{o_{2}^{\prime}}^{\downarrow}+(U-3J)\hat{n}_{o_{1}^{\prime}}^{s^{\prime}}\hat{n}_{o_{2}^{\prime}}^{s^{\prime}}-J\hat{c}^{{\dagger}}_{o_{1}^{\prime},\uparrow}\hat{c}_{o_{1}^{\prime},\downarrow}\hat{c}^{{\dagger}}_{o_{2}^{\prime},\downarrow}\hat{c}_{o_{2}^{\prime},\uparrow}+J\hat{c}^{{\dagger}}_{o_{1}^{\prime},\uparrow}\hat{c}^{{\dagger}}_{o_{1}^{\prime},\downarrow}\hat{c}_{o_{2}^{\prime},\downarrow}\hat{c}_{o_{2}^{\prime},\uparrow}\right)\,.$ (56)
Here $U$ and $J$ are measured in units of $t$ which is set to $1$. We consider weak coupling where the TPSC formalism should give the correct leading order behavior. For the ED calculations, we use pyED [107] from the TRIQS framework [108]. The results for all three methods as a function of $U$ and $J$ are plotted in Fig. 2.
Both TPSC3 and TPSC5 show indeed the correct leading order behavior for all TPEVs, irrespective of the ratio of $U$ and $J$. Further, we observe that TPSC5 performs significantly worse when $J$ becomes unphysically large, while TPSC3 stays much closer to the ED result. This trend also persists at lower temperatures, see Fig. 7 in Appendix B. The TPEV do match well between TPSC3 and ED throughout the whole range of examined parameters. In general, the deviation is strongest in the pair-hopping component, in which the error monotonously increases with $J$. We stress again that since we stay in the weak-coupling regime, these benchmarks have no direct implications for applications in the strongly correlated regime.
III.3.2 Comparison to DMFT
@@
As a next step, we compare the results from TPSC to DMFT. For this, we consider the same model as Ref. [88]
$\hat{H}=\sum_{s,\braket{i,j},o}tc^{\dagger}_{o,s,i}c^{\dagger}_{o,s,j}+\hat{H}_{HK}\,,$ (57)
where $t=1$ is fixed and the inverse temperature is $\beta=2$. For completeness, we include the results from the implementation proposed in Ref. [88] (here named TPSC). The DMFT calculations are performed using w2dynamics [64].
@@
This model can be seen as a worst-case scenario as in the absence of inter-orbital hopping the off-diagonal occupations are guaranteed to be zero. Hence, the $P$-channel as well as opposite spin $C$-channel (in the case of TPSC5) Ansatz equations are ill-defined, making a renormalization of the corresponding vertices impossible. In other words, the corresponding Hartree-Fock decouplings always result in a net-zero contribution of these terms to the self-energy. Interestingly, even though the Ansatz equations might break down, the derivation remains valid – we can still renormalize these components, albeit in a less controlled fashion. We found that fixing the renormalization of the channels to that of the one channel in which the Ansatz is not ill-defined works for a wide range of parameters.
@@
Both proposed TPSC approaches show promising results in the weak-coupling regime, but the deviation from DMFT rapidly increases with stronger interactions. Assuming that the DMFT results are accurate, the error is largest in the absence of Hund’s coupling and becomes smaller for larger $J$, as expected from our discussion above. In general TPSC5 seems to outperform TPSC3, even though both fulfill the internal consistency check in the small to intermediate coupling region. Notably, there is no significant improvement in the case of large Hund’s couplings when compared to the approach put forward in Ref. [88].
III.4 Comparison of interacting susceptibilities
In addition to the quality of the TPEV’s we can also test the interacting charge and spin susceptibilities. In the following we will compare our results to D-TRILEX [103], again for a model of the form of Fig. 1, however introducing a hopping imbalance between the red and blue orbitals
$H=\sum_{\alpha,\braket{i,j},s}t_{\alpha}c^{\dagger}_{\alpha,s,i}c_{\alpha,s,j}+H_{HK}$ (58)
with $t_{\mathrm{red}}=1$ and $t_{\mathrm{blue}}=0.75$. $U$ and $J$ are in the following given in units of $t_{\mathrm{red}}$.
@@ III.4 Comparison of interacting susceptibilities
Figure 5: Comparison of the charge (upper row) and spin (lower row) susceptibility between TPSC3 (blue), TPSC5 (orange), TPSC from Ref. [66] (green) and D-TRILEX (purple) at different $U$ and $J=0.2U$ at $\beta=2$.
We find that both TPSC algorithms over/underestimate the spin/charge susceptibility in comparison to D-TRILEX, which was also found in Ref [93]. However, the overestimation is weaker for TPSC5 and stronger for TPSC3. Most notably, the charge susceptibility is even stronger suppressed than in prior TPSC formulations.
IV Conclusion and Outlook
@@
In this paper we introduced two variants of a fully self-consistent multi-orbital TPSC approach dubbed TPSC3 and TPSC5. We analyzed the structure of the equations and showed analytically that in the limit of vanishing Hund’s coupling the Ansatz equations result in the same TPEVs as for the single orbital model, which is unphysical. Further, we provided an explanation for this behavior in terms of the equivalence between the Hatree-Fock decomposition and a minimization in terms of single Slater determinants, highlighting that representing influences from many low energy states onto the TPEVs are not fully reproducible by a single state with modified prefactors. Furthermore, by considering the local Hilbert space we provided an understanding of why the approach performs better at larger $J/U$. Notably, this understanding allows to assess the expected quality of the results from TPSC by inspecting the local spectrum. On a qualitative level, we found that the TPEVs from TPSC5 outperform the ones from TPSC3 when compared to DMFT. Furthermore, in the limiting case of $J=U/3$ the new variants do not outperform the conceptually simpler but approximate approach suggested in Ref. [88].
@@
While TPSC itself cannot be improved much without fundamentally modifying the basic approach, the recently proposed combination of TPSC and DMFT [93, 89] resolves partially the issue of wrong TPEVs due to the Hartree-Fock decoupling but also ensures the correct handling of correlations at the zeroth-order level. Thus, in multi-orbital systems, the application of TPSC in combination DMFT appears as the most promising route to correct for the flaws uncovered in the present work and in Ref. [89]. A major advantage of TPSC in this formulation is that no vertex needs to be extracted from the DMFT simulation, which offers a numerically much cheaper alternative to D-TRILEX and related approaches [14, 19].
V Acknowledgements
@@ Appendix A Kadanoff-Baym Formalism
In the following we derive the central equations for the multi-orbital TPSC approach, following Ref. [78, 88]. Our starting point is the Green’s function generating functional [109]
$\mathcal{G}[\phi]=-\ln\langle T_{\tau}e^{-c^{\dagger}_{1^{\prime}}c_{2^{\prime}}\phi_{1^{\prime},2^{\prime}}}\rangle=-\ln\langle T_{\tau}S[\phi]\rangle=\ln(Z[\phi])\,.$ (59)
@@
We use numbers as a short hand for a collection of spin, orbital, position and time indices. Primed variables are summed over. The external source field $\phi$ has been introduced as a mathematical trick and at the end of the calulation it can be set to 00. The single particle Green’s function is obtained as the functional derivative of $\mathcal{G}$ with respect to $\phi$
$G_{1,2;\phi}=\frac{\delta\mathcal{G}}{\delta\phi(2,1)}=-\frac{\braket{T_{\tau}S[\phi]c_{1}c^{\dagger}_{2}}}{Z[\phi]}\,.$ (60)
@@
With this, we perform a Legendre transformation giving us the vertex generating functional or Luttinger-Ward functional $\Phi[G]$ whose first derivative is the self-energy
$\Sigma_{1,2;\phi}=-\frac{\delta\Phi[G]}{\delta G(2,1)}\,.$ (61)
From the Legendre transform relation between the generating functionals the Dyson-equation follows as
$G^{-1}=G_{0}^{-1}-\phi-\Sigma\,.$ (62)
The second functional derivative of $\mathcal{G}[\phi]$ is the generalized susceptibility given as
$\chi_{1423}=\frac{\delta G_{1,2;\phi}}{\delta\phi_{3,4}}=\braket{c_{1}c^{\dagger}_{2}c^{\dagger}_{3}c_{4}}_{\phi}+\braket{c_{1}c^{\dagger}_{2}}_{\phi}\braket{c_{4}c^{\dagger}_{3}}_{\phi}\,,$ (63)
while the vertex can be expressed as
$\Gamma_{1,2,3,4}=\frac{\delta\Sigma_{1,2}}{\delta G_{3,4}}\,.$ (64)
With these we obtain a Dyson-equation like connection between the interacting and non-interacting susceptibility:
00 0 $=\frac{\delta(G^{-1}_{1,3^{\prime}}G_{3^{\prime},3})}{\delta{\phi_{4,2}}}$ (65)
$=G^{-1}_{1,3^{\prime}}\frac{\delta(G_{3^{\prime},3})}{\delta{\phi_{4,2}}}+\frac{\delta(G^{-1;0}_{1,3^{\prime}}-\phi_{1,3^{\prime}}-\Sigma_{1,3^{\prime}})}{\delta{\phi_{4,2}}}G_{3^{\prime},3}$ (66)
$\Leftrightarrow\chi_{1,2,3,4}$ $=-G_{1,1^{\prime}}\frac{(\delta\phi_{1^{\prime},3^{\prime}}+\Sigma_{1^{\prime},3^{\prime}})}{\delta{\phi_{4,2}}}G_{3^{\prime},3}$ (67)
$=\chi^{0}_{1,2,3,4}-G_{1,1^{\prime}}\frac{\delta\Sigma_{1^{\prime},3^{\prime}}}{\delta{G_{4^{\prime},2^{\prime}}}}\frac{\delta G_{4^{\prime},2^{\prime}}}{\delta{\phi_{4,2}}}G_{3^{\prime},3}$ (68)
$=\chi^{0}_{1,2,3,4}+\chi^{0}_{1,3^{\prime},3,1^{\prime}}\Gamma_{1^{\prime},3^{\prime},4^{\prime},2^{\prime}}\chi_{4^{\prime},2,2^{\prime},4}\,,$ (69)
which is the general form of the Bethe-Salpeter equation. In a $SU(2)$ symmetric system, we can further use that all occuring quantities are diagonal in physical spin-space, thus resulting in spin-channel-specific equations which read (pulling out spin indices from the numbers)
$\chi^{S^{i^{\prime}},S^{j^{\prime}}}_{1,2,3,4}$ $=(\chi^{0}_{1,2,3,4}+\chi^{0}_{1,3^{\prime},3,1^{\prime}}\Gamma_{1^{\prime},3^{\prime},4^{\prime},2^{\prime}}\chi_{4^{\prime},2,2^{\prime},4})\sigma^{i}_{s_{1},s_{3}}\sigma^{j}_{s_{2},s_{4}}$ (70)
$\Leftrightarrow\chi_{1,2,3,4}^{S^{z},S^{z}}$ $=\chi_{1,2,3,4}^{0;S^{z},S^{z}}+\chi^{0;S^{z},S^{z}}_{1,3^{\prime},3,1^{\prime}}\Gamma_{1^{\prime},3^{\prime},4^{\prime},2^{\prime}}^{S^{z},S^{z}}\chi_{4^{\prime},2,2^{\prime},4}^{S^{z},S^{z}}$ (71)
$\Leftrightarrow\chi_{1,2,3,4}^{S^{0},S^{0}}$ $=\chi_{1,2,3,4}^{0;S^{0},S^{0}}-\chi^{0;S^{0},S^{0}}_{1,3^{\prime},3,1^{\prime}}\Gamma_{1^{\prime},3^{\prime},4^{\prime},2^{\prime}}^{S^{0},S^{0}}\chi_{4^{\prime},2,2^{\prime},4}^{S^{0},S^{0}}\,,$ (72)
where $S_{x}S_{x}$, $S_{y}S_{y}$ and $S_{z}S_{z}$ are related by $SU(2)$ symmetry.
The equation of motion is a central building block for TPSC, and for completeness we derive it here for a Hamiltonian with local and static two-electron interactions,
$H=t_{13}c^{\dagger}_{3}(\tau)c_{1}(\tau)+\frac{1}{2}U_{1234}c^{\dagger}_{3}(\tau)c^{\dagger}_{4}(\tau)c_{2}(\tau)c_{1}(\tau)\,.$ (73)
@@
In this case, the Heisenberg equation of motion reads
$\partial_{\tau_{5}}c_{5}(\tau_{5})$ $=\left[H,c_{5}(\tau_{5})\right]$ (74)
$=\left[t_{1^{\prime}3^{\prime}}c^{\dagger}_{3^{\prime}}(\tau)c_{1^{\prime}}(\tau),c_{5}(\tau_{5})\right]+\frac{1}{2}\left[U_{1^{\prime}2^{\prime}3^{\prime}4^{\prime}}c^{\dagger}_{3^{\prime}}(\tau)c^{\dagger}_{4^{\prime}}(\tau)c_{2^{\prime}}(\tau)c_{1^{\prime}}(\tau),c_{5}(\tau_{5})\right]$ (75)
$=-t_{1^{\prime}3^{\prime}}\{c_{3^{\prime}}^{\dagger},c_{5}(\tau_{5})\}c_{1^{\prime}}+\frac{1}{2}U_{1^{\prime}2^{\prime}3^{\prime}4^{\prime}}\left(c_{3^{\prime}}^{\dagger}(\tau)c_{4^{\prime}}^{\dagger}(\tau)\{c_{2^{\prime}}(\tau)c_{1^{\prime}}(\tau),c_{5}(\tau_{5})\}-\{c_{3^{\prime}}^{\dagger}(\tau)c_{4^{\prime}}^{\dagger}(\tau),c_{5}(\tau_{5})\}c_{2^{\prime}}(\tau)c_{1^{\prime}}(\tau)\right)$ (76)

### v3

@@ 3.1 Model with pure density-density interactions
Let us first answer the second question. In the single-orbital case we fix the local and static expectation values to be the exact ones as determined by the sum rules. Crucially, the single site contains no further internal structure meaning that the exact local expectation value is always proportional to the non-interacting one. Thus representing this expectation value locally by a renormalized Hartree-Fock expectation value works, i.e., we can always get the correct value of the double occupancy by multiplying the non-interacting double occupancy with a single number. Multi-orbital systems have additional degrees of freedom, which enter the local and static contribution. We still approximate the TPEVs in Eq. (8) by a Hartree-Fock decoupling. In other words, we approximate the ground state by a single Slater-determinant. However, the local Hilbert-space is now spanned by more than a single wave-function and all of these states contribute to the expectation value. Thus, introducing interactions can fundamentally alter the structure of the TPEV’s such that these can deviate from a behavior multiplicably representable by the single Slater determinant from Hartree-Fock, which explains the issue we described in the first question posed above. As an example for this, let us consider a single dimer (i.e. a single site in Fig. 1) at half-filling with a Hubbard-Kanamori interaction. For simplicity, let us call the two orbitals $a$ and $b$. The Hamiltonian then contains only interaction terms and reads
$H=\sum_{i\in\{a,b\}}U\hat{n}_{o_{i}}^{\uparrow}\hat{n}_{o_{i}}^{\downarrow}+$ $\left((U-2J)\hat{n}_{a}^{\uparrow}\hat{n}_{b}^{\downarrow}+(U-3J)\hat{n}_{a}^{s^{\prime}}\hat{n}_{b}^{s^{\prime}}\right.$ (59)
$\quad\left.-J\hat{c}^{{\dagger}}_{a,\uparrow}\hat{c}_{a,\downarrow}\hat{c}^{{\dagger}}_{b,\downarrow}\hat{c}_{b,\uparrow}+J\hat{c}^{{\dagger}}_{a,\uparrow}\hat{c}^{{\dagger}}_{a,\downarrow}\hat{c}_{b,\downarrow}\hat{c}_{b,\uparrow}\right)\,.$
This we can readily diagonalize in the subspace containing two electrons (|$\ket{\uparrow_{a},\downarrow_{a}}$, |$\ket{\uparrow_{a},\downarrow_{b}}$, |$\ket{\uparrow_{b},\downarrow_{a}}$, |$\ket{\uparrow_{a},\uparrow_{b}}$, |$\ket{\downarrow_{a},\downarrow_{b}}$, |$\ket{\uparrow_{b},\downarrow_{b}}$). We find three different groups of eigenstates: The first one is three-fold degenerate with eigenvalue $U-3J$ and spanned by |$\ket{\uparrow_{a},\uparrow_{b}}$, |$\ket{\downarrow_{a},\downarrow_{b}}$ and (|$(\ket{\uparrow_{a},\downarrow_{b}}-\ket{\uparrow_{b},\downarrow_{a}})/\sqrt{2}$. These three states form the $S=1$ sector. Next, we do have two states with eigenvalue $U-J$, (|$(\ket{\uparrow_{a},\downarrow_{b}}+\ket{\uparrow_{b},\downarrow_{a}})/\sqrt{2}$ and (|$(\ket{\uparrow_{a},\downarrow_{a}}-\ket{\uparrow_{b},\downarrow_{b}})/\sqrt{2}$. And lastly, a single eigenstate with eigenvalue $U+J$, (|$(\ket{\uparrow_{a},\downarrow_{a}}+\ket{\uparrow_{b},\downarrow_{b}})/\sqrt{2}$.
Therefore, in the absence of $J$ all six eigenvalues are degenerate in the local problem. In Hartree-Fock (and by inheritance in TPSC) only one of these six states is represented by the decoupling and the others are missed entirely (thus behaving like the single orbital model in the absence of $J$). Thereby, the basic idea of TPSC, which relies on fixing the local and static expectation values starting from Hartree-Fock, breaks down due to the inadequateness of Hartree-Fock. Furthermore, since this issue seems to be inherited from the Hartree-Fock decoupling, which is central to TPSC, there is no simple repair for this problem without fundamentally altering the structure of the method.
3.2 Analysis of the model with $J=U/3$
If $J=U/3$, the same-spin density-density interaction becomes exactly zero. This choice is interesting, as it is the limiting case where the approach in Ref. [90] and TPSC5 show the best agreement compared to DMFT, see Fig. 3. What is the origin of this apparent improvement?
First, we should note that this is not a magic value at which the results suddenly agree, but there appears to be a steady improvement with increasing $J$. Second, we observe that for $J=U/3$ the same-spin interaction is exactly 00 and stays at that value throughout the calculation in TPSC5. In TPSC3 this interaction will gain a nonzero value, leading to the instability we observe in Fig. 3.
Since we found in the last section that the problems arising can be traced back to the Hartree-Fock ansatz, the question arises whether we can again understand the improvement as an inherited effect. To test this, we again consider the half-filled dimer described above. The three groups of eigenvalues we found above split with increasing $J$. Thus, while the point $U=3J$ is not special, the separation between the excited state and the ground state sectors grows linearly with $J$. Furthermore, the $S=1$ subspace is special, as two states only contribute to a single TPEV ($\braket{n^{\uparrow}_{0}n^{\uparrow}_{1}}$) and the other state essentially behaves like the single orbital model. Therefore, one should be able to approximate the TPEV’s by a scaling in all components but the $\braket{n^{\uparrow}_{0}n^{\uparrow}_{1}}$ one, which is incorrectly estimated because of the missing degenerate states. This is exactly what is observed for both TPSC5 and in Ref. [90]. Thus we conclude that in the end this improvement is related to the structure and number of states determining the local and static expectation values. Furthermore, the improved performance is entirely explained by the improved performance of the Hartree-Fock ansatz.
From these observations, we can extract a guiding principle of when to expect TPSC to work reliably in a multi-orbital setting and when not: If there is a gap in the local spectrum between the ground state and the excited states, TPSC should perform better than if the states are close in energy. Furthermore, when analyzing the nature of the states, we can extract which TPEV is expected to deviate strongly and which not. This is intrinsically tied to Hartree-Fock being essentially a ground state targeting approach which performs better at lower temperatures. However, the local and static approximation of the vertex becomes more and more inappropriate at lower temperatures which is why TPSC typically fails at low temperatures.
In summary, the regimes in which multi-orbital TPSC is guaranteed to work well are the weak coupling limit and a small region in which excited states do not contribute to the local TPEVs (controlled by temperature and the gap), while the vertex still is approximately local and static. The former criterion is accessible by inspection, while the latter one is not clear a priori and requires other numerical simulations to be fully gauged. Notably, the issue described above is partially resolved when starting from DMFT TPEVs [91, 95], since fixing the double occupancies through an external source prevents the negative feedback from a wrong starting point.
3.3 Numerical benchmarks
@@
Keeping the above described shortcomings in mind, we are now exploring the range of applicability of multi-orbital TPSC by testing the reliability of TPSC3 and TPSC5 through comparison to more accurate numerical approaches. In a first step, we exclusively aim at validating the quality of the predicted TPEVs, as these form the basis for the TPSC approach. For this we compare to both DMFT and D-Trilex. It should be stressed that while neither of the approaches are exact, both methods have proven to be suitable for the considered parameter regime [50, 19].
3.3.1 Comparison to DMFT - TPEV’s
@@
First, we compare the results from TPSC to DMFT. For this, we consider the same model as Ref. [90], which we already introduced above, see Eq. (46). The inverse temperature is set to $\beta=2$, thus no sharp features in momentum space appear. We run the simulations on a $24\times 24$ momentum mesh and utilize the sparse-ir [108, 109, 110] to compress the single particle Green’s function with $\omega_{\text{max}}=30$ and a tolerance of $10^{-12}$ for the singular values. For completeness, we also include the results from the implementation proposed in Ref. [90] (here named TPSC). The DMFT calculations are performed using w2dynamics [64].
@@
The considered model can be seen as a worst-case scenario as in the absence of inter-orbital hopping the off-diagonal occupations are guaranteed to be zero. Hence, the $P$-channel as well as the opposite spin $C$-channel (in the case of TPSC5) Ansatz equations are ill-defined, making a renormalization of the corresponding vertices impossible. In other words, the corresponding Hartree-Fock decouplings always results in a net-zero contribution of these terms to the self-energy. Interestingly, even though the Ansatz equations might break down, the derivation remains valid – we can still renormalize these components, albeit in a less controlled fashion. We found that fixing the renormalization of the channels to that of the one channel in which the Ansatz is not ill-defined works for a wide range of parameters.
@@
Both proposed TPSC approaches show promising results in the weak-coupling regime, but the deviation from DMFT rapidly increases with stronger interactions. The error is largest in the absence of Hund’s coupling and becomes smaller for larger $J$, as expected from our discussion above. In general TPSC5 seems to outperform TPSC3, even though both fulfill the internal consistency check in the small to intermediate coupling region. Notably, there is no significant improvement in the case of large Hund’s couplings when compared to the approach put forward in Ref. [90].
@@
Figure 4: Comparison of different self-energy components at the first Matsubara frequency between TPSC3 (blue), TPSC5 (orange), TPSC from Ref. [90] (green) and D-TRILEX (purple) at different $U$ indicated by different background colors and $J=0.25U$. The inverse temperature is fixed to $\beta=2$.
Figure 5: Comparison of the charge (upper row) and spin (lower row) susceptibility between TPSC3 (blue), TPSC5 (orange), TPSC from Ref. [90] (green) and D-TRILEX (purple) at different $U$ and $J=0.25U$ at $\beta=2$.
3.3.2 Comparison to D-TRILEX - self-energy and susceptibilities
In addition to the quality of the TPEV’s we can also asses the quality of interacting charge and spin susceptibilities as well as the self-energy. In the following we will compare our results to D-TRILEX results provided by the authors of Ref. [19]. For this, we consider a model of the form of Fig. 1 where an additional hopping imbalance is introduced between the red and blue orbitals
$H=-\sum_{\alpha,\braket{i,j},s}t_{\alpha}c^{\dagger}_{\alpha,s,i}c_{\alpha,s,j}+H_{HK}\;,$ (60)
with $t_{\mathrm{red}}=1$ and $t_{\mathrm{blue}}=0.75$. $U$ is in the following given in units of $t_{\mathrm{red}}$, while $J$ is fixed to $0.25U$. To obtain smooth curves along the irreducible path without the need for interpolation we enlarge the momentum mesh to $120\times 120$. Furthermore, we compare quantities obtained without analytical continuation to rule it out as a source of error.
First, let us consider the self-energy at the lowest Matsubara frequency (Fig. 4). Qualitatively, all TPSC variants agree with D-TRILEX at both interactions values considered. Quantitatively, the TPSC results are not fully agreeing with each other which is some what expected from the different formulations. We note that TPSC3 appears to be furthest away from the D-TRILEX self-energy for this specific model, while TPSC5 and Ref. [66] appear to be better agreeing.
A similar picture emerges for the susceptibility: all TPSC variants over(under)-estimate the spin(charge) susceptibility in comparison to D-TRILEX, in agreement with what was reported in Ref. [95]. However, the overestimation is weaker for TPSC5 and stronger for TPSC3. Most notably, the charge susceptibility is even more strongly suppressed than in prior TPSC formulations. This disagreement between the TPSC variants is a consequence of the vertex having a different from, see App. B.1, which leads to problems in the absence of $J$, see Ref. [90] but better agreement to the reference data at large $J$, see Fig. 3.
4 Conclusions and Outlook
@@
In this paper we introduced two variants of a fully self-consistent multi-orbital TPSC approach dubbed TPSC3 and TPSC5. We analyzed the structure of the equations and showed analytically that in the limit of vanishing Hund’s coupling the Ansatz equations result in the same TPEVs as for the single orbital model, which is unphysical. Furthermore, we provided an explanation for this behavior in terms of the equivalence between the Hatree-Fock decomposition and a minimization in terms of single Slater determinants, highlighting that the influences from many low-energy states onto the local TPEVs cannot be fully reproduced by a single state with modified prefactors. By considering the local Hilbert space we also provided an understanding of why the approach performs better at larger $J/U$. Notably, this understanding allows to assess the expected quality of the results from TPSC by inspecting the local spectrum. On a qualitative level, we found that the TPEVs from TPSC5 outperform the ones from TPSC3 when compared to DMFT. In the limiting case of $J=U/3$ the new variants do not outperform the conceptually simpler but approximate approach suggested in Ref. [90].
@@
While TPSC itself cannot be improved much without fundamentally modifying the basic approach, the recently proposed combination of TPSC and DMFT [95, 91] not only partially resolves the issue of wrong TPEVs due to the Hartree-Fock decoupling but also ensures the correct handling of correlations at the zeroth-order level. Thus, in multi-orbital systems, the application of TPSC in combination with DMFT appears as the most promising route to correct for the flaws uncovered in the present work and in Ref. [91]. A major advantage of TPSC in this formulation is that no vertex needs to be extracted from the DMFT simulation, which offers a numerically much cheaper alternative to D-TRILEX and related approaches [14, 19].
5 Acknowledgements
@@ Appendix A Kadanoff-Baym Formalism
In the following we derive the central equations for the multi-orbital TPSC approach, following Ref. [80, 90]. We will only consider objects with the same number of in- and out-going legs. Therefore, we will stick to the definition that the first half of the indices are always in-going legs, while the second half are outgoing legs. Furthermore, to keep the equations more managable we will use numbers as a short hand for a collection of quantum numbers $1\equiv(o_{1},s_{1},\boldsymbol{r}_{1},\tau_{1})$. A kronecker-delta in numbers is defined as the product of the Kronecker deltas/Delta-distributions of the individual quantum numbers.
Our starting point is the Green’s function generating functional [111]
$\mathcal{G}[\phi]=-\ln\left(\Braket{T_{\tau}e^{-c^{\dagger}_{o_{3^{\prime}}s_{3^{\prime}}}(\boldsymbol{\tau}_{3^{\prime}})c_{o_{1^{\prime}}s_{1^{\prime}}}(\boldsymbol{\tau}_{1^{\prime}})\phi_{o_{1^{\prime}}s_{1^{\prime}},o_{3^{\prime}}s_{3^{\prime}}}(\boldsymbol{\tau}_{1^{\prime}},\boldsymbol{\tau}_{3^{\prime}})}}_{S}\right)=-\ln\left(\Braket{T_{\tau}S[\phi]}_{S}\right)=\ln(Z[\phi])\,,$ (61)
@@
where the expectation value over the sources is calculated with regard to the action of the full system $S$. Primed variables are summed over. The external source field $\phi$ has been introduced as a mathematical trick and at the end of the calculation it should be set to 00. The single-particle Green’s function is obtained as the functional derivative of $\mathcal{G}$ with respect to $\phi$,
$G_{o_{1},o_{3};\phi}^{s_{1},s_{3}}(\boldsymbol{\tau}_{1},\boldsymbol{\tau}_{3})$ $=\frac{\delta\mathcal{G}[\phi]}{\delta\phi_{o_{1},o_{3}}^{s_{1},s_{3}}(\boldsymbol{\tau}_{3},\boldsymbol{\tau}_{1})}=-\frac{\braket{T_{\tau}S[\phi]c_{o_{1},s_{1}}(\boldsymbol{\tau}_{1})c^{\dagger}_{o_{3},s_{3}}(\boldsymbol{\tau}_{3})}_{S}}{Z[\phi]}$ (62)
$=-\braket{T_{\tau}c_{o_{1},s_{1}}(\boldsymbol{\tau}_{1})c^{\dagger}_{o_{3},s_{3}}(\boldsymbol{\tau}_{3})}_{\phi}\,.$
@@
The subscript $\phi$ indicates that the expectation value is to be calculated with the source fields added. With this, we perform a Legendre transformation resulting in the vertex generating functional or Luttinger-Ward functional $\Phi[G]$. The first functional derivative of $\Phi[G]$ is the self-energy
$\Sigma_{o_{1},o_{3};\phi}^{s_{1},s_{3}}(\boldsymbol{\tau}_{1},\boldsymbol{\tau}_{3})=-\frac{\delta\Phi[G]}{\delta G_{o_{3},o_{1};\phi}^{s_{3},s_{1}}(\boldsymbol{\tau}_{3},\boldsymbol{\tau}_{1})}\,.$ (63)
In the next step, we derive the equation of motion and through that obtain the Dyson equation relating the dressed Green’s function and the self-energy. Our starting point is a general Hamiltonian with local and static interactions
$H=t_{o_{1},o_{3}}^{s_{1},s_{3}}(\boldsymbol{r}_{1},\boldsymbol{r}_{3})c^{\dagger}_{o_{3},s_{3}}(\boldsymbol{r}_{3})c_{o_{1},s_{1}}(\boldsymbol{r}_{1})+\frac{1}{2}U_{o_{1},o_{2},o_{3},o_{4}}^{s_{1},s_{2},s_{3},s_{4}}c^{\dagger}_{o_{3},s_{3}}(\boldsymbol{r})c^{\dagger}_{o_{4},s_{4}}(\boldsymbol{r})c_{o_{2},s_{2}}(\boldsymbol{r})c_{o_{1},s_{1}}(\boldsymbol{r})\,.$ (64)
@@
In this case, the Heisenberg equation of motion (in imaginary time) reads
$\partial_{\tau_{5}}c_{o_{5},s_{5}}(\boldsymbol{r}_{5},\tau_{5})$ $=\left[H,c_{o_{5},s_{5}}(\boldsymbol{r}_{5},\tau_{5})\right]$ (65)
$=t_{o_{1^{\prime}},o_{3^{\prime}}}^{s_{1^{\prime}},s_{3^{\prime}}}(\boldsymbol{r}_{1}^{\prime},\boldsymbol{r}_{3}^{\prime})\left[c^{\dagger}_{o_{3^{\prime}},s_{3^{\prime}}}(\boldsymbol{r}_{3}^{\prime})c_{o_{1^{\prime}},s_{1^{\prime}}}(\boldsymbol{r}_{1}^{\prime}),c_{o_{5},s_{5}}(\boldsymbol{r}_{5})\right]_{H}$ (66)
$\quad+\frac{1}{2}U_{o_{1^{\prime}},o_{2^{\prime}},o_{3^{\prime}},o_{4^{\prime}}}^{s_{1^{\prime}},s_{2^{\prime}},s_{3^{\prime}},s_{4^{\prime}}}\left[c^{\dagger}_{o_{3},s_{3}}(\boldsymbol{r}^{\prime})c^{\dagger}_{o_{4},s_{4}}(\boldsymbol{r}^{\prime})c_{o_{2},s_{2}}(\boldsymbol{r}^{\prime})c_{o_{1},s_{1}}(\boldsymbol{r}^{\prime}),c_{o_{5},s_{5}}(\boldsymbol{r}_{5})\right]_{H}$
$=-t_{o_{1^{\prime}},o_{3^{\prime}}}^{s_{1^{\prime}},s_{3^{\prime}}}(\boldsymbol{r}_{1}^{\prime},\boldsymbol{r}_{3}^{\prime})\left(\{c^{\dagger}_{o_{3^{\prime}},s_{3^{\prime}}}(\boldsymbol{r}_{3}^{\prime}),c_{o_{5},s_{5}}(\boldsymbol{r}_{5})\}c_{o_{1^{\prime}},s_{1^{\prime}}}(\boldsymbol{r}_{1}^{\prime})\right)_{H}$ (67)
$\quad+\frac{1}{2}U_{o_{1^{\prime}},o_{2^{\prime}},o_{3^{\prime}},o_{4^{\prime}}}^{s_{1^{\prime}},s_{2^{\prime}},s_{3^{\prime}},s_{4^{\prime}}}\left([c^{\dagger}_{o_{3^{\prime}},s_{3^{\prime}}}(\boldsymbol{r}^{\prime})c^{\dagger}_{o_{4^{\prime}},s_{4^{\prime}}}(\boldsymbol{r}^{\prime}),c_{o_{5},s_{5}}(\boldsymbol{r}_{5})]c_{o_{2^{\prime}},s_{2^{\prime}}}(\boldsymbol{r}^{\prime})c_{o_{1^{\prime}},s_{1^{\prime}}}(\boldsymbol{r}^{\prime})\right)_{H}$
$=-t_{o_{1^{\prime}},o_{3^{\prime}}}^{s_{1^{\prime}},s_{3^{\prime}}}(\boldsymbol{r}_{1}^{\prime},\boldsymbol{r}_{3}^{\prime})c_{o_{1^{\prime}},s_{1^{\prime}}}(\boldsymbol{r}_{1}^{\prime},\tau_{5})\delta_{o_{3^{\prime}},o_{5}}\delta_{s_{3^{\prime}},s_{5}}\delta_{\boldsymbol{r}_{3}^{\prime},\boldsymbol{r}_{5}}$ (68)